# zkFuzz: Foundation and Framework for Effective Fuzzing of Zero-Knowledge Circuits

### v1

@@ 7. Related Work
Dynamic Testing. Fuzzing and mutation testing have been used to detect bugs in ZKP infrastructures (Hochrainer et al., 2024; Leeb, 2024; Xiao et al., 2025), typically focusing on compiler bugs. In contrast, our approach targets security vulnerabilities in individual ZK circuits.
@@
Table 6. Comparison between zkFuzz and existing detection methods. A partial circle represents partial capability. zkFuzz demonstrates superior performance across all criteria, offering comprehensive automatic testing capabilities.
@@
Circomspect (of Bits, 2024) Yes
Picus (Pailoor et al., 2023) No
ConsCS (Jiang et al., 2025) No
AC4 (Chen et al., 2024) No
ZKAP (Wen et al., 2024) Yes
@@
8. Limitations and Future Work
Limitations. zkFuzz adopts a fuzzing approach for vulnerability detection and thus inherits certain limitations. For example, the execution time of zkFuzz is proportional to that of the target program, and slicing and partially executing the target might improve zkFuzz. In addition, our implementation currently supports only Circom, and potential improvements include covering other DSLs.
Future Work. Integrating zkFuzz with formal verification, static analyzers, and ML/LLM-based bug detection (Zhang et al., 2025; She et al., 2019; He et al., 2024; Gan et al., 2024) might further boost its scalability and flexibility. Additionally, zkFuzz could serve as a subroutine for fuzzing ZK compilers like Cairo (Goldberg et al., 2021), Noir (Aztec Network, 2024), and Leo (Chin et al., 2021), helping to detect discrepancies between compiled programs and circuit constraints.
9. Conclusion
We introduce zkFuzz, a fuzzing framework for detecting ZK circuit bugs. It uses TCCT, a comprehensive and language-independent definition of ZK circuit bugs. zkFuzz applies program mutation and target-guided artificial inputs to find under- and over-constrained circuits without false positives. It discovered 38 previously unknown bugs in real-world circuits.
Appendix A Proofs
Proof of Theorem 3.9.
$\mathcal{C}=\phi$
$\mathcal{P}(x)=\begin{cases}((),(1))&\quad\text{if all clauses in$\phi$ is satisfied with $x$}\\ (\bot,\bot)&\quad\text{otherwise}\end{cases}$
$\mathcal{C}=false$
(Complexity of TCCT): Next, we prove the Trace-Constraint Consistency Test’s co-NP-completeness by reducing the Boolean tautology problem, a known co-NP-complete problem, to it.
Lemma A.1.
Proof.
Lemma A.2.
Proof.
$\phi_{1}$ is a tautology if and only if $\phi_{1}\equiv\hbox{TRUE}$
Given two Boolean formulas $\phi_{1}$ and $\phi_{2}$, we construct an instance of the Trace-Constraint Consistency Test $\langle\mathcal{P},\mathcal{C}\rangle$ over $\mathbb{F}_{2}$ as follows:
$\mathcal{P}((x_{1},\ldots,x_{n}))=\begin{cases}((),(1))&\text{if $(x_{1},\ldots,x_{n})$ satisfies $\phi_{1}$}\\ ((),(0))&\text{otherwise}\\ \end{cases}$
$\mathcal{C}(x,z,y)=\begin{cases}1&\text{if $x$ satisfies $\phi_{2}$ and $y=1$}\\ 1&\text{if $x$ does not satisfy $\phi_{2}$ and $y=0$}\\ 0&\text{otherwise}\end{cases}$
This reduction can be performed in polynomial time. Since the Boolean tautology problem is co-NP-complete (Arora and Barak, 2009) and we have shown a polynomial-time reduction to the Trace-Constraint Consistency Test, we conclude that it is co-NP-complete.
Appendix B Additional Examples of ZK Program
Correct Implementation of RShift1. Code 9 presents a secure implementation of a 1-bit right shift in Circom. The input is first converted into a bit array, the shift operation is applied to this array, and the result is then converted back into the output.
1template RShift1(N) {
2 signal input x;
3 signal output y;
4
5 component x_bits = Num2Bits(N);
6 x_bits.in <== x;
7
8 signal y_bits[ N - 1 ];
9 for (var i = 0; i < N - 1; i++) {
10 y_bits[i] <== x_bits.out[i + 1];
11 }
12
13 component y_num = Bits2Num(N - 1);
14 y_num.in <== y_bits;
15 y <== y_num.out;
16}
Code 9: Secure implementation of 1-bit right shift.
LessThan. While explicit constraints are readily apparent in circuit designs, implicit assumptions used in Circom’s de facto standard library, circomlib (iden3, 2025), can also introduce subtle under-constrained bugs if not properly accounted for.
Consider the ZK program in Code 11 from zk-regex (zkemail, 2025) which validates the range of the given characters and is expected to abnormally terminate when they contain values larger than 255: This template uses the LessThan circuit from circomlib, which implicitly assumes that both inputs are represented by n bits or fewer. Although this assumption is well known(iden3 contributors, 2023; of Bits, 2024), the EmailAddrRegex template does not validate the bitwidth of the inputs provided to LessThan.
The critical vulnerability arises from the potential overflow in Line 7 and 8 of LessThan when input a is excessively large. Note that Num2Bits is a template converting an input to the bit expression. To illustrate, suppose $q=401$ and msg[0] $=400$. The computation in Line 7 of LessThan becomes $400+(1\ll 8)-255\equiv 0\ \bmod 401$. Hence, LessThan outputs $1$, erroneously passing the range check of EmailAddrRegex. Our fuzzer successfully discovered this bug, which the developer subsequently confirmed.
1template LessThan(n) {
2 assert(n <= 252);
3 signal input in[2];
4 signal output out;
5 component n2b = Num2Bits(n+1);
6
7 n2b.in <== in[0] + (1<<n) - in[1];
8 out <== 1 - n2b.out[n];
9}
Code 10: Implementation of LessThan template from circomlib (iden3, 2025). This template assumes that the inputs are at most n bits wide. If the inputs exceed n bits, the comparison result may be incorrect.
1template EmailAddrRegex(n) {
2 signal input msg[n];
3
4 signal in_range_checks[n];
5 for (var i = 0; i < n; i++) {
6 in_range_checks[i] <== LessThan(8)([msg[i], 255]);
7 in_range_checks[i] $=$ 1;
8 // the rest is omitted
Code 11: Example of an under-constrained circuit from zk-regex (zkemail, 2025). LessThan implicitly assumes that the bit length of both inputs are not longer than n, although LessThan itself does not check the range of inputs.
1template SplitReward() {
2 signal input x;
3 signal z;
4 signal output y;
5
6 z <-- x \ 2;
7 z * 2 $=$ x;
8 y <== z + 1;
9}
Code 12: Example of over-constrained circuit. Suppose assert is not added in the computation for $=$.
Table 7. Trace and Constraint Satisfaction Sets of Over-Constrained Circuit ($q=5$).
$\mathcal{T}(\mathcal{P})$ $\mathcal{S}(\mathcal{C})$
x z y x z y
0 0 1 0 0 1
1 0 1 1 3 4
2 1 2 2 1 2
3 1 2 3 4 0
4 2 3 4 2 3
[…]
Zero-Division Pattern. Code 13 presents a real-world example of an under-constrained circuit from circomlib (iden3, 2025) that exhibits the zero-division vulnerability pattern. On line 13, a division is performed where the numerator is a degree-2 polynomial in in[0], while the denominator is a degree-1 polynomial in in[1].
1template MontgomeryDouble() {
2 signal input in[2];
3 signal output out[2];
4
5 var a = 168700;
6 var d = 168696;
7 var A = (2 * (a + d)) / (a - d);
8 var B = 4 / (a - d);
9
10 signal lamda;
11 signal x1_2;
12
13 x1_2 <== in[0] * in[0];
14
15 lamda <-- (3*x1_2 + 2*A*in[0] + 1 ) / (2*B*in[1]);
16 lamda * (2*B*in[1]) $=$ (3*x1_2 + 2*A*in[0] + 1 );
17
18 out[0] <== B*lamda*lamda - A - 2*in[0];
19 out[1] <== lamda * (in[0] - out[0]) - in[1];
20}
Code 13: Real example of Zero Division pattern in circomlib (iden3, 2025). lambda can be any value when both the denominator and the numerator of Line 15 are 0, making this circuit non-deterministically under-constrained.
Hash-Check Pattern. Code 14 illustrates the real-world circuit implementing the hash-check pattern from the maci project (Explorations, 2020). Although this circuit is under-constrained due to insecure call of LessThan within Quin TreeInclusionProof, detecting this issue without the hash-check is difficult, as finding a valid input and Merkle root pair requires inverting the hash function.
1template QuinLeafExists(levels){
2 var LEAVES_PER_NODE = 5;
3 var LEAVES_PER_PATH_LEVEL = LEAVES_PER_NODE - 1;
4 var i;
5 var j;
6
7 signal input leaf;
8 signal input path_elements[levels][LEAVES_PER_PATH_LEVEL];
9 signal input path_index[levels];
10 signal input root;
11
12 // Verify the Merkle path
13 component verifier = QuinTreeInclusionProof(levels);
14 verifier.leaf <== leaf;
15 for (i = 0; i < levels; i ++) {
16 verifier.path_index[i] <== path_index[i];
17 for (j = 0; j < LEAVES_PER_PATH_LEVEL; j ++) {
18 verifier.path_elements[i][j] <== path_elements[i][j];
19 }
20 }
21
22 root $=$ verifier.root;
23}
Code 14: Real example of Hash Check pattern from maci (Explorations, 2020). Determining a valid pair of input and Merkle root requires inverting the hash function. The circuit is under-constrained due to an unsafe invocation of LessThan within QuinTreeInclusionProof.
Appendix C Details of Benchmark Datasets
Figure 6. Distribution of circuit characteristics across 354 real-world ZKP benchmarks. The dataset includes circuits with large specifications, underscoring the need for scalable bug-detection tools. Notably, approximately 60% of circuits contain weak assignments, posing a potential risk of under-constrained vulnerabilities in real-world circuits.
Fig. 6 presents a histogram of the number of instructions in the program $\mathcal{P}$ after loop-unfolding, the constraint length $|\mathcal{C}|$, and the number of weak assignments in the dataset. The x-axis is logarithmic. We observe that nearly 60% of the benchmarks contain weak assignments, which may lead to under-constrained bugs.
Figure 7. Top 10 Frequently Occurring Circomlib Templates.
Fig.7 also presents the top 10 most frequently used templates in our benchmark, all of which originate from circomlib(iden3, 2025). The two most commonly used templates are Num2Bits and IsZero. Notably, IsEqual internally relies on IsZero, while LessThan builds upon Num2Bits. Templates ranked 5 through 9—namely Sigma, MixLast, Ark, and MixS—are mainly used within Poseidon (Grassi et al., 2021), a widely adopted hash function in zero-knowledge (ZK) circuits.
Appendix D Additional Experiments and Analysis
D.1. RQ6: Generality of TCCT
We investigate how our TCCT framework is generalized compared to existing formulations.
Consider Code 15, which presents a lightly revised version of Code 4. Recall that assert does not add a constraint, leading to an under-constrained circuit. Line 4 and 9 are added by us to prevent any signal from being unused in the constraint, while the circuit is still under-constrained. Our TCCT can cover this issue since the constraints accept trace even if b $\neq$ h.y, and zkFuzz successfully detects this bug. In contrast, none of the existing tools can detect this bug as the constraints are deterministic, and all signals are utilized in the constraints.
1template Verify() {
2 signal input a;
3 signal input b;
4 signal output c;
5
6 component h = Hash();
7 h.x <== a;
8 assert(b==h.y);
9 c <== h.y * b;
10}
Code 15: Example of an under-constrained circuit undetectable by other existing tools. Only our TCCT model is capable of capturing this bug, and zkFuzz successfully identifies it.
Another example is Code 16, a real-world circuit from the iden3 protocol, where the GreaterThan internally uses LessThan, and the IN does not verify the bit-length of GreaterThan’s inputs. Consequently, ZKAP erroneously marks the circuit as unsafe. However, the first argument to GreaterThan is count—the sum of outputs from IsEqual. Since each IsEqual outputs either 0 or 1, the maximum value of count is bounded by valueArraySize (which is 3 in this case). As the second argument is 0, both inputs to GreaterThan remain sufficiently small, ensuring the circuit is safe. Notably, zkFuzz does not generate any false positive for this circuit.
1template IN (valueArraySize){
2 signal input in;
3 signal input value[valueArraySize];
4 signal output out;
5
6 component eq[valueArraySize];
7 var count = 0;
8 for (var i=0; i<valueArraySize; i++) {
9 eq[i] = IsEqual();
10 eq[i].in[0] <== in;
11 eq[i].in[1] <== value[i];
12 count += eq[i].out;
13 }
14 // GreaterThan internally uses LessThan
15 component gt = GreaterThan(252);
16 gt.in[0] <== count;
17 gt.in[1] <== 0;
18 out <== gt.out;
19}
20
21component main = IN(3);
Code 16: Real-world circuit from iden3: This is safe despite misleading lack of bit-length checks
Result 6: Generality of TCCT TCCT is more general than existing models, as it can cover bugs that others miss while avoiding false positives in real-world circuits.
D.2. RQ7: Alternative Mutation Strategies
Figure 8. Various input generation and program mutation strategies.
Result 7: Alternative Mutation Strategies Adding two additional mutation strategies, random constant addition and expression deletion, does not significantly improve the performance of zkFuzz, highlighting the effectiveness of its default mutation strategies.
D.3. RQ8: Performance of Other Baselines
Tab. 8 presents the performance of the modified Circomspect, which recursively analyzes all internally called templates, and ZKAP with the Unconstrained-Signal detector. While recursive Circomspect detects all vulnerabilities, it also increases false positives, causing precision to drop sharply from 0.50 to 0.37. For ZKAP, the Unconstrained-Signal detector does not contribute to finding new bugs in our benchmarks.
Table 8. Performance of Circomspect with recursive analysis and ZKAP with the US and the USCO detector. Prec. stands for precision.
Constraint Size
Circomspect
ZKAP
(with US and USCO)
TP FP Prec. TP FP Prec.
Small 41 20 0.67 33 20 0.62
Medium 6 20 0.23 1 23 0.04
Large 16 27 0.37 3 5 0.38
Very Large 6 46 0.12 2 14 0.13
Total 69 113 0.38 39 62 0.39
Result 8: Performance of Other Baselines The original ZKAP and Circomspect with recursive analysis detect more bugs, but at the cost of significantly increasing false positives and reducing precision.
D.4. RQ9: Over-Constrained Circuits
To validate zkFuzz’s ability to detect over-constrained circuits, we conduct experiments with the constraint_ assert_ dissabled flag enabled, allowing over-constrained circuits to exist. In this mode, the equality constraint $=$ adds a condition to the constraints set without inserting an assert into the program. Consequently, developers likely need to add the corresponding assert manually in the Circom file to prevent over-constrained circuits. Since most circuits are designed to function under the Circom compiler’s default configuration, enabling this flag renders many circuits over-constrained. In our benchmark, zkFuzz identifies 159 such cases out of 354 test cases. However, because these circuits typically run under the default compiler settings, such issues are unlikely to pose a security risk in practice.
Result 9: Over-Constrained Circuits Since most circuits are designed for the default settings of the Circom compiler, enabling the constraint_ assert_ dissabled flag causes nearly half of the circuits to be classified as over-constrained.

### v2

@@ 7 Related Work
Dynamic Testing. Fuzzing and mutation testing have been used to detect bugs in ZKP systems [26, 27, 28], typically focusing on compiler bugs. In contrast, we target security vulnerabilities in individual ZK circuits.
8 Limitations and Future Work
Limitations. zkFuzz adopts a fuzzing approach for bug detection and thus inherits certain limitations. For example, the execution time is proportional to the size of the target program, and slicing and partially executing the target might improve zkFuzz. Our prototype for Noir also remains preliminary, lacking guiding and target selection methods.
Future Work. A primary avenue for improvement is to enhance Noir support by adding more guiding and target selectors. In addition, since TCCT does not depend on any particular DSL and many DSLs compile down to standardized constraint formats such as R1CS, TCCT - and consequently zkFuzz - can be applied across a wide range of ZK languages by targeting these common trace/constraint representations, regardless of surface-level differences in syntax or semantics. Integrating zkFuzz with formal verification, static analyzers, and ML-based bug detection [59, 60, 61, 62] might also further boost its scalability and flexibility.
9 Conclusion
We present zkFuzz, a fuzzing framework for finding ZK circuit bugs based on TCCT —a language-independent bug model. zkFuzz uses guided mutations and inputs to detect under- and over-constrained circuits without false positives, uncovering 59 previously unknown bugs in real-world circuits.
Appendix A Proofs
Proof of Theorem 1.
$\mathcal{C}=\phi$
[…]
$\mathcal{C}=false$
(Complexity of TCCT): Next, we prove TCCTs’s co-NP-completeness by reducing the Boolean tautology problem, a known co-NP-complete problem, to it.
Lemma 2.
Proof.
Lemma 3.
Proof.
$\phi_{1}$ is a tautology if and only if $\phi_{1}\equiv\hbox{TRUE}$
Given two Boolean formulas $\phi_{1}$ and $\phi_{2}$, we construct an instance of TCCT $\langle\mathcal{P},\mathcal{C}\rangle$ over $\mathbb{F}_{2}$ as follows:
[…]
This reduction can be performed in polynomial time. Since the Boolean tautology problem is co-NP-complete [63] and we have shown a polynomial-time reduction to TCCT, we conclude that it is co-NP-complete.
Appendix B Additional Examples of ZK Program
Correct Implementation of RShift1. Code 10 presents a secure implementation of a 1-bit right shift in Circom. The input is first converted into a bit array, the shift operation is applied to this array, and the result is then converted back into the output.
1template RShift1(N) {
2 signal input x;
3 signal output y;
4 signal out <== Num2Bits(N)(x);
5 signal bits[N-1];
6 for(var i=0; i<N-1; i++) {bits[i] <== out[i+1];}
7 y <=== Bits2Num(N - 1)(bits); }
Code 10: Secure implementation of 1-bit right shift.
LessThan. While explicit constraints are readily apparent in circuit designs, implicit assumptions used in Circom’s de facto standard library, circomlib [44], can also introduce subtle under-constrained bugs if not properly accounted for. For example, LessThan circuit from circomlib, which checks whether the first input is lower than the second input, implicitly assumes that both inputs are represented by n bits or fewer. The critical vulnerability arises from the potential overflow in Line 7 of LessThan when input a is excessively large. This is due to the modular arithmetic operations over $\bmod q$. To illustrate, consider a case where $q=17$, $a=10$, and $b=1$. The computation in Line 7 of LessThan becomes $10+(1\ll 3)-1\equiv 0\ \bmod 17$, erroneously indicating that $a<b$.
This class of vulnerabilities, stemming from the violation of implicit assumptions of templates, can be incorporated into our TCCT framework. The easiest way for this case is to insert one additional assert to the computation $\mathcal{P}$: assert((in[0] < in[1] and out == 1) or (in[0] >= in[1] and out == 0)).
Consider the real-world ZK program in Code 12 from zk-regex [31] which validates the range of the given characters and is expected to abnormally terminate when they contain values larger than 255: This template uses the LessThan circuit, although the EmailAddrRegex template does not validate the bitwidth of the inputs provided to LessThan. To illustrate the attack vector, suppose $q=401$ and msg[0] =400=400. The computation in Line 7 of LessThan becomes $400+(1\ll 8)-255\equiv 0\ \bmod 401$. Hence, LessThan outputs $1$, erroneously passing the range check of EmailAddrRegex. Our fuzzer successfully discovered this bug, which the developer subsequently confirmed.
1template LessThan(n){
2 assert(n <= 252);
3 signal input in[2];
4 signal output out;
5 component n2b = Num2Bits(n+1);
6 n2b.in <== in[0] + (1<<n) - in[1];
7 out <== 1 - n2b.out[n];}
Code 11: LessThan template from circomlib [44]. This template assumes that the inputs are at most n bits wide. If the inputs exceed n bits, the comparison result may be incorrect.
1template EmailAddrRegex(n){
2 signal input msg[n];
3
4 signal in_range_checks[n];
5 for (var i=0; i<n; i++) {
6 in_range_checks[i] <== LessThan(8)([msg[i], 255]);
7 in_range_checks[i] === 1;
8 // the rest is omitted
Code 12: Example of an under-constrained circuit from zk-regex [31]. LessThan implicitly assumes that the bit length of both inputs are not longer than n, although LessThan itself does not constrain the range of inputs.
1template SplitReward(){
2 signal input x;
3 signal z;
4 signal output y;
5 z <-- x \ 2;
6 z * 2 === x;
7 y <== z + 1;}
Code 13: Example of over-constrained circuit. Suppose assert is not added in the computation for ===.
TABLE VI: Trace and Constraint Satisfaction Sets of Code 13 ($q=5$).
$\mathcal{T}(\mathcal{P})$ $\mathcal{S}(\mathcal{C})$
x z y x z y
0 0 1 0 0 1
1 0 1 1 3 4
2 1 2 2 1 2
3 1 2 3 4 0
4 2 3 4 2 3
However, enabling the constraint _assert _disabled flag can lead to over-constrained circuits. Code 13 presents a modified version of the Reward from [13], showing a ZK program that performs integer division. Note that $\backslash$” operator in Circom is the integer division, while ”*” operator is the multiplication modulo $q$. Suppose $q=5$ and constraint_assert_disabled is turned on. Then, Tab. VI shows the trace and constraint satisfaction sets. $\mathcal{S}(\mathcal{C})$ does not contain $(\text{x}=1,\text{z}=0,\text{y}=1)$ and $(\text{x}=3,\text{z}=1,\text{y}=2)$ of $\mathcal{T}(\mathcal{P})$, meaning that it is over-constrained ($\mathcal{T}(\mathcal{P})\setminus\mathcal{S}(\mathcal{C})\neq\emptyset$). Note that it is also under-constrained.
Hash-Check Pattern. Code 14 illustrates the real-world circuit implementing the hash-check pattern from the maci project [64]. Although this circuit is under-constrained due to insecure call of LessThan within Quin TreeInclusionProof, detecting this issue without the hash-check is difficult, as finding a valid input and Merkle root pair requires inverting the hash function.
1template QuinLeafExists(levels){
2 var LEAVES_PER_NODE = 5;
3 var LEAVES_PER_PATH_LEVEL = LEAVES_PER_NODE - 1;
4 var i, j;
5
6 signal input leaf, root;
7 signal input path_index[levels];
8 signal input path_elements[levels][LEAVES_PER_PATH_LEVEL];
9
10 // Verify the Merkle path
11 component verifier = QuinTreeInclusionProof(levels);
12 verifier.leaf <== leaf;
13 for (i = 0; i < levels; i ++) {
14 verifier.path_index[i] <== path_index[i];
15 for (j = 0; j < LEAVES_PER_PATH_LEVEL; j ++) {
16 verifier.path_elements[i][j] <== path_elements[i][j];}}
17 root === verifier.root;
18}
Code 14: Real example of Hash Check pattern from maci [64]. Determining a valid pair of input and Merkle root requires inverting the hash function.
Appendix C Non-Deterministic Witness Generation
TCCT’s generality allows it to avoid false positives that arise in prior definitions like Picus [18], Ac4 [24], and Conscs [25] when dealing with inherently non-deterministic programs. While currently widely used DSLs like Circom are deterministic, it is plausible that future DSLs like CODA [21] might introduce inherent non-determinism (e.g., through concurrency). In such a setting, a program $\mathcal{P}$ could legitimately produce multiple valid outputs y, with constraints $\mathcal{C}$ designed to accept exactly those outputs. In this case, TCCT correctly classifies $(\mathcal{P},\mathcal{C})$ as well-constrained, while prior works will produce a false positive. To understand the effect, consider the following Circom-like program inspired by the real Decoder circuit in circomlib. Here, random_binomial non-deterministically returns either 0 or 1. The Decoder takes input inp and produces two outputs:
  • out, a binary array decoded from inp, and
  • success, a boolean flag indicating whether decoding succeeds (specifically, decoding is successful if and only if the inp-th entry of out is 1).
In this case, although parts of the output (out) are non-deterministic, the constraints ensure validity only when the values remain consistent with the program’s definition. Under TCCT, such a program is therefore not classified as under-constrained. In contrast, prior formulations that consider only the constraints—without accounting for the program semantics—would incorrectly label this circuit as under-constrained, leading to a false positive.
1template Decoder(w) {
2 signal input inp;
3 signal output out[w], success;
4 Var lc = 0;
5
6 for (var i = 0; i < w; i++) {
7 out[i] <-- random_binomial();
8 out[i] * (inp - i) === 0;
9 lc = lc + out[i];
10 }
11 lc ==> success;
12 success * (success - 1) === 0;
13}
Code 15: Example of non-deterministic but well-constrained ZK program. TCCT correctly concludes that this is well-constrained.
Appendix D Additional Experiments and Analysis
RQ7: Alternative Mutation Strategies
Figure 6: Various input generation and program mutation strategies.
D.1 RQ8: Performance of Other Baselines
Tab. VII presents the performance of the modified Circomspect, which recursively analyzes all internally called templates, and ZKAP with the Unconstrained-Signal detector. While recursive Circomspect detects all vulnerabilities, it also increases false positives, causing precision to drop sharply to 0.33. As discussed in § 3, Circomspect cannot detect misuse of assert. However, in this dataset, it happens to identify all circuits containing real bugs, mainly because it incorrectly flags code that is not actually problematic. For ZKAP, the Unconstrained-Signal detector does not contribute to finding new bugs in our benchmarks.
TABLE VII: Performance of Circomspect with recursive analysis and ZKAP with the US and the USCO detector.
Constraint Size
Circomspect
ZKAP
(with US and USCO)
TP FP Prec. TP FP Prec.
Small 58 28 0.67 37 23 0.61
Medium 7 23 0.23 2 29 0.06
Large 17 30 0.36 4 8 0.33
Very Large 6 100 0.05 2 29 0.06
Total 88 181 0.33 45 89 0.34
D.2 RQ9: Impact Assessment of Vulnerabilities
Table VIII presents an analysis of the impacts of previously unknown bugs discovered by zkFuzz, categorizing each unique bug into seven types: (1) incorrect primitive operations, such as faulty integer division or string matching; (2) game cheating, such as falsifying the winner; (3) identity forgery, such as bypassing age verification; (4) cryptographic computation bugs, such as incorrect hash or ciphertext computations; (5) ownership forgery, such as maliciously claiming possession of specific data; (6) incorrect reward calculation, such as errors in incentive mechanisms; and (7) others. The most common type is ownership forgery, mainly due to incorrect range checks and misuse of the assert statement.
TABLE VIII: Bug impact categorization in previously unknown ZK bugs identified by zkFuzz.
Type #Bugs
Incorrect Primitive Operation 11
Game Cheating 7
Identity Forgery 8
Cryptographic Computation Bug 4
Ownership Forgery 15
Incorrect Reward Calculation 7
Others 7
D.3 RQ10: Over-Constrained Circuits
To validate zkFuzz’s ability to detect over-constrained circuits, we conduct experiments with the constraint_ assert_ disabled flag enabled, allowing over-constrained circuits to exist. In this mode, the equality constraint === adds a condition to the constraints set without inserting an assert into the program. Consequently, developers likely need to add the corresponding assert manually in the Circom file to prevent over-constrained circuits. Since most circuits are designed to function under the Circom compiler’s default configuration, enabling this flag renders many circuits over-constrained. In our benchmark, zkFuzz identifies 258 such cases out of 452 test cases. However, because these circuits typically run under the default compiler settings, such issues are unlikely to pose a security risk in practice.
Appendix E Comparison with Existing Works
Tab. IX presents a comparison between the capabilities of existing works and our zkFuzz along different dimensions, demonstrating its superiority.
@@ Appendix E Comparison with Existing Works
TABLE IX: Comparison between zkFuzz and existing detection methods. A partial circle represents partial capability. zkFuzz demonstrates superior performance across all criteria, offering comprehensive automatic testing capabilities.
@@
Circomspect [20] Yes
Picus [18] No
ConsCS [25] No
AC4 [24] No
ZKAP [13] Yes